OpenStudy (anonymous):

Determine the sum of each infinite geometric series, if it exists. a) t1=8, r=-1/4 b)t1=3, r=4/3

OpenStudy (anonymous):

let, the first term=t1 and common ratio=r so the sum of the infinite series, =t1/(1-r) for r<1 =t1/(r-1) for r>1 now you evaluate for the two cases.

OpenStudy (anonymous):

yup, i got for a) Sinfinity=6 1/2, and for b) Sinfinity=-9

OpenStudy (anonymous):

but the back of the book says that b) has no sum... how do you know if it has a sum or not?

OpenStudy (anonymous):

oh sorry. I made a mistake. there will be no sum for infinite series with r>1 only sum for infinite series, t1/(1-r) for |r|<1

OpenStudy (anonymous):

oookkkayy..... but in b) r is more than 1....(r=4/3)

OpenStudy (anonymous):

yes. actually the sum for the second case diverges as the series is divergent. so we can nor calculate the infinite sum.

OpenStudy (anonymous):

i think the back of the book is wrong... cause wouldn't it only have no sum if r is 0

OpenStudy (anonymous):

if r=0 then the series will have only the first term.

OpenStudy (anonymous):

the series is, a, ar, ar^2, ar^3, ar^4,..............................